Find a linear function g such that
g(9) =-21 and g(7)=-15

Answer by stanbon(75887)

Find a linear function g such that
g(9) =-21 and g(7)=-15
----
You have 2 solutions: (9,-21) and (7,-15)
------
slope = (-15--21)/(7-9) = 6/-2 = -3
-----
Form: y = mx + b
Solve for "b":
-15 = -3*7 + b
---
b = 6
--------
Equation:
y = -3x + 6
===============
Cheers,
Stan H.
================